# Membership Inference on Text-to-image Diffusion Models via Conditional Likelihood Discrepancy

### v1

@@ Abstract
Text-to-image diffusion models have achieved tremendous success in the field of controllable image generation, while also coming along with issues of privacy leakage and data copyrights. Membership inference arises in these contexts as a potential auditing method for detecting unauthorized data usage. While some efforts have been made on diffusion models, they are not applicable to text-to-image diffusion models due to the high computation overhead and enhanced generalization capabilities. In this paper, we first identify a conditional overfitting phenomenon in text-to-image diffusion models, indicating that these models tend to overfit the conditional distribution of images given the text rather than the marginal distribution of images. Based on this observation, we derive an analytical indicator, namely Conditional Likelihood Discrepancy (CLiD), to perform membership inference. This indicator reduces the stochasticity in estimating the memorization of individual samples. Experimental results demonstrate that our method significantly outperforms previous methods across various data distributions and scales. Additionally, our method shows superior resistance to overfitting mitigation strategies such as early stopping and data augmentation.
@@ 1 Introduction
We conduct extensive experiments on three text-to-image datasets [29, 57, 27] with different distributions and sizes using the mainstream open-sourced text-to-image diffusion models [8, 41] under both fine-tuning and pre-training settings. First, our methods consistently outperform existing baselines across various data distributions and training scenarios. For example, $\text{CLiD}_{vec}$ achieves average improvements over the best baseline by 28.7%, 29.8%, and 55.9% in terms of ASR, AUC, and TPR@1%FPR metrics, respectively. Second, we suggest evaluating the membership inference methods by observing the effectiveness trajectory with increasing training steps of a model. The experimental results show our methods reveal membership earlier with fewer training steps. Next, ablation studies further demonstrate the effect of our CLiD indicator, even with fewer query count, our method still outperforms baseline methods (Fig. 3). Last, experiments show that our methods exhibits stronger resilience to data augmentation, and resilience to even adaptive defenses.
@@ 3.2 Conditional Overfitting Phenomenon
We further compute the change in FID after truncating the condition and observe that the change in FID of the member set is consistently greater than that of the hold-out set (Fig. 1 (b)), which inspires revealing membership by this overfitting discrepancy. Recalling the aim of text-to-image diffusion model is to fit a latent space mapping from text to image, image data augmentation is commonly used to enhance the model generalization. For instance, the official fine-tuning script of Hugging-Face [19] employs RandCrop and Randomflip as the default augmentation [20]. However, few trainers disturb the text condition as it is discrete and such disturbance would result in a decline of model utility (Sec. 4.5). Therefore, we believe that leveraging this phenomenon contributes to addressing the challenges of the strong generalization of text-to-image diffusion models with the resistance to data augmentation.
@@ 3.4 Implementation of CLiD-MI
Vector-based attack–$\text{CLiD}_{vec}$. We combine the estimated values of Eq. Eq. (14) and Eq. (15) to obtain the feature vectors corresponding to each data point:
@@ 3.5 Practical Considerations
Classifiers of $\text{CLiD}_{vec}$. Due to the simplicity of the feature vectors, we do not need a neural network as the classifier [11]. Simpler classifiers help to prevent overfitting. In our experiments, we utilize XGBoost [6] and utilize the predict confidence.
@@ 4.1 Setups
Training setups. Previous membership inference on text-to-image diffusion models usually relies on strong overfitting settings. To evaluate the performance more realistically, we consider the two following setups for fine-tuning: (1) Over-training. Following the previous work [11, 23, 13], we fine-tune 15,000 steps on Pokemon datasets, and 150,000 steps on MS-COCO and Flickr (with only $/2500$ 2500 dataset size). (2) Real-world training. Considering that trainers typically do not train for such high steps, we recalibrate the steps based on the training steps/dataset size ratio (approximately 20) of official fine-tuning scripts on Huggingface [19]. Thus, we train 7,500 steps, 50,000 steps and 200,000 steps for the Pokémon, MS-COCO and Flickr datasets, respectively. Additionally, we employ the default data augmentation in training codes [20] to simulate real-world scenarios.
@@
Evaluation metrics. We follow the widely used metrics of previous works [3, 4, 11, 23, 13], including ASR (i.e., the accuracy of membership inference), AUC and the True Positive Rate (TPR) when the False Positive Rate (FPR) is 1% (i.e. TPR@1%FPR).
@@ 4.3 Performance on Various Training Steps
From Tab. 1 and Tab. 2, we find that the training steps greatly influence the effectiveness of membership inference. All membership inference methods tend to exhibit satisfactory performance when the model is trained for an excessive number of steps that conflict with real-world scenarios. Therefore, we emphasize that the effectiveness trajectory of membership inference across varying training steps should also be utilized to evaluate different methods. Better membership inference methods should reveal membership information earlier as training progresses.
@@ 4.4 Ablation Study
Figure 3: Performance of $\text{CLiD}_{th}$ under various Monte Carlo sampling count. The legend labels are sorted in ascending order of AUC values and SecMI. $M/N=0$ indicates that the corresponding term is not utilized in Eq. (16).
@@
Monte Carlo sampling numbers. In Fig. 3, we observe that when setting $M=N$, the performance improves as the number of Monte Carlo sampling increases. And the performance is improved slightly when $M,N>3$. Hence, we set $M,N=3$ to ensure the balance between a low query number and satisfied performance. Moreover, the results of "M=1, N=1" and "SecMI" also demonstrate that $\text{CLiD}_{th}$ outperforms previous works even with fewer query counts.
@@ 4.5 Resistance to Defense
Since data augmentation is commonly used in training and can mitigate the effectiveness of membership inference, we use it to evaluate the performance of methods under defense. As the baseline methods already exhibit weak performance under real-world training setup, we opt not to incorporate additional data augmentation. Instead, we remove the default data augmentation from training scripts [20] to observe the effectiveness change of different methods. We fine-tune Stable Diffusion models for 50,000 steps with MS-COCO, report the metrics, and calculate the metrics changes in Tab. 3. We observe that the effectiveness of all membership inference methods declines after data augmentation is introduced during training. Note that PFAMI [13] exhibits the highest sensitivity to data augmentation since it infers membership by probability fluctuation after images are perturbed, which also explains its significant performance decline between Tab. 1 and Tab. 2. Compared to the baselines, our method exhibits the smallest decrease, which indicates the stronger resilience of our methods to data augmentation.
Adaptive defense. We further consider adaptive defense: assuming the trainers is aware of our methods and perturb the text of image-text datasets before training. We consider the following adaptive methods: (1) rephrasing the original text  [^3] We utilize ChatGPT-3.5 with the following prompt: ”Please rewrite the following sentences while keeping the key semantics.”, (2) randomly deleting 10%, 30% words in text, and (3) shuffling 50% of the image-text mappings in the dataset. In Tab. 5, we observe that except for shuffling, the other adaptive defense methods have almost no effect on $\text{CLiD}_{th}$. And shuffling damages the model utility (too high FID values), rendering this defense meaningless.
@@
ASR AUC FPR@1%FPR ASR ($\Delta$) AUC ($\Delta$) FPR@1%FPR ($\Delta$)
@@ 4.6 Performance on Pre-trained Models
To ensure the distribution consistency between the member and hold-out set, we respectively select $2500$ samples from the LAION-Aesthetics v2 5+ and LAION-2B MultiTranslated [45] as member/hold-out set following [12], and conduct membership inference on Stable Diffusion v1-5 [40]. As shown in Tab. 5, our method consistently outperforms the baselines across all three metrics.
@@ 5 Related Works
Membership inference on diffusion models. Recently, several membership inference methods designed for diffusion models have been proposed. In the grey-box or white-box setting, Carlini et al. [4] firstly conduct membership inference on unconditional diffusion models by conducting LiRA (Likelihood Ratio Attack) [3], with the requirement of training multiple shadow models. Matsumoto et al. [32] make the first step by utilizing diffusion loss to conduct query-based membership inference. Some works [11, 23] leverage the DDIM [48] deterministic forward process [22] to access the posterior estimation errors of diffusion models. And Fu et al. [13] leverage the probability fluctuations by perturbing image samples. Few works consider the black-box settings [54, 35]. However, these studies either assume partial knowledge of member set data [54] or assume extensive fine-tuning steps ($100\sim 500$ epochs) [35], both of which do not align with real-world scenarios.
@@ 6 Conclusion
To extensively validate the effectiveness of Assumption 3.1, we utilize additional metrics as the distances metric $D$ in Eq. (8), including Wasserstein Distance [55], Kernel MMD (Maximum Mean Discrepancy) [55] and the 1-Nearest Neighbor Classifier (1-NN) [31], in addition to FID [15]. As observed in Fig. A.1, regardless of the metric used for $D$, Assumption 3.1 consistently holds, thereby confirming the broad applicability of Conditional Overfitting phenomenon.
@@ D.1 Monte Carlo Sampling
In Fig. D.1, we observe that the single Monte Carlo estimation achieves optimal accuracy when $t_{i}\in[400,500]$. Similar results are shown in [28]. Therefore, consistent with [28], we sample at intervals of $10$ centered around the timestep $450$. For example, if the estimation number (i.e., $M,N$ in Eq (14)) is $3$, we use the time list of $[440,450,460]$. This difference may arise from the different signal-to-noise ratios of images with various resolution at the same diffusion timestep [18]. Note that [4, 11] indicate that for DDPM of Cifar-10 [25] the best estimation timestep is around $100$. This finding suggests that different Monte Carlo sampling times should be designed for diffusion models of different scales.
@@ D.2 Reduction Methods
In implementation, we actually reduce the condition $\mathbf{c}$ diversely and calculate $p_{\theta}(\mathbf{x}|\mathbf{c})$ to approximate $p_{\theta}(\mathbf{x})$. In this part, we evaluate the effectiveness of different reduction methods. We consider three methods in Sec. 3.5: (1) Simply Clipping. We simply use the first, middle, and last thirds of the sentences as text inputs. (2) Gaussian Noise. We add Gaussian noises with the scales of $50\%,70\%,90\%$ to the overall text embeddings. (3) Importance Clipping. We calculate the importance of words in the text [^4] https://github.com/ma-labo/PromptCharm [49, 51] and replace them with “pad” tokens in descending order by varying proportions of $30\%,50\%,70\%$. For all three methods, we additionally use the null text as a $\mathbf{c}^{*}$. The experiments are conducted on the real-world training setting with MS-COCO dataset. And we also employ null text solely to compute Eq. (14) without reduction methods for comparison.
@@ Appendix E Compute Overhead and Resources
Computational Overhead. As a query-based member inference method, the computational efficiency of our method primarily depends on the number of queries. A lower number of queries signifies a more efficient member inference method. Our method significantly outperforms the baselines when the numbers of queries are about the same (such as SecMI and PFAMI in Sec. 4.2). Furthermore, even with a fewer number of queries, our method exhibits a noticeable improvement compared to the baselines (Sec. 4.4).
Compute Resources. Our experiments are divided into two main parts: training (fine-tuning) and inference, both conducted on a single RTX A6000 GPU. The time of execution in the training phase depends on the training steps. For example, we perform $7,500$, $50,000$, and $200,000$ steps for Pokemon [27], MS-COCO [29] and Flickr [57] dataset, which take about 2 hours, 12 hours, and 48 hours, respectively. The time of execution in inference time depends on the methods’ query numbers. For example, with the query number of $15$, our membership inference method on a dataset of size $/2500$ 2500 takes approximately $80$ minutes per run for all data points. Typically, we perform this inference once on the shadow model and once on the target model, resulting in a total time cost of $160$ minutes.
@@ Appendix G Limitation
Despite the significant improvements in membership inference of text-to-image diffusion models on various data distributions and data sizes, this work still has limitations. Firstly, due to the limited availability of open-sourced pre-trained weights of text-to-image diffusion models, evaluations under the pre-training setting are not sufficiently comprehensive. Secondly, the Conditional Overfitting phenomenon discussed in this paper is only explored within the domain of text-to-image diffusion models. From an insight perspective, its presence may not be such limited. Future efforts could explore this in conditional diffusion models across different modalities such as text-to-video, and further investigate more universally applicable membership inference methods.

### v2

@@ Abstract
Text-to-image diffusion models have achieved tremendous success in the field of controllable image generation, while also coming along with issues of privacy leakage and data copyrights. Membership inference arises in these contexts as a potential auditing method for detecting unauthorized data usage. While some efforts have been made on diffusion models, they are not applicable to text-to-image diffusion models due to the high computation overhead and enhanced generalization capabilities. In this paper, we first identify a conditional overfitting phenomenon in text-to-image diffusion models, indicating that these models tend to overfit the conditional distribution of images given the corresponding text rather than the marginal distribution of images only. Based on the observation, we derive an analytical indicator, namely Conditional Likelihood Discrepancy (CLiD), to perform membership inference, which reduces the stochasticity in estimating memorization of individual samples. Experimental results demonstrate that our method significantly outperforms previous methods across various data distributions and scales. Additionally, our method shows superior resistance to overfitting mitigation strategies such as early stopping and data augmentation.
@@ 1 Introduction
We conduct extensive experiments on three text-to-image datasets [29, 57, 27] with different distributions and sizes using the mainstream open-sourced text-to-image diffusion models [8, 41] under both fine-tuning and pre-training settings. First, our methods consistently outperform existing baselines across various data distributions and training scenarios. For example, $\text{CLiD}_{vec}$ achieves average improvements over the best baseline by 28.7%, 29.8%, and 55.9% in terms of ASR, AUC, and TPR@1%FPR metrics, respectively. Second, we suggest evaluating the membership inference methods by observing the effectiveness trajectory with increasing training steps of a model. The experimental results show our methods reveal membership earlier with fewer training steps. Next, ablation studies further demonstrate the effect of our CLiD indicator, even with fewer query count, our method still outperforms baseline methods (Fig. 3). Last, experiments show that our methods exhibit stronger resistance to data augmentation, and exhibit resistance to even adaptive defenses.
@@ 3.2 Conditional Overfitting Phenomenon
We further compute the change in FID after truncating the condition and observe that the change in FID of the member set is consistently greater than that of the hold-out set (Fig. 1 (b)), which inspires revealing membership by this overfitting discrepancy. Recalling the aim of text-to-image diffusion model is to fit a latent space mapping from text to image, image data augmentation is commonly used to enhance the model generalization. For instance, the official fine-tuning script of Hugging-Face [19] employs Random-Crop and Random-Flip as the default augmentation [20]. However, few trainers disturb the text condition as it is discrete and such disturbance would result in a decline of model utility (Sec. 4.5). Therefore, we believe that leveraging this phenomenon contributes to addressing the challenges of the strong generalization of text-to-image diffusion models with the resistance to data augmentation.
@@ 3.4 Implementation of CLiD-MI
Vector-based attack–$\text{CLiD}_{vec}$. We combine the estimated values of Eq. (14) and Eq. (15) to obtain the feature vectors corresponding to each data point:
@@ 3.5 Practical Considerations
Classifiers of $\text{CLiD}_{vec}$. Due to the simplicity of the feature vectors, we do not need a neural network as the classifier [11]. Simpler classifiers help to prevent overfitting. In our experiments, we utilize XGBoost [6] and utilize its predict confidence.
@@ 4.1 Setups
Training setups. Previous membership inference on text-to-image diffusion models usually relies on strong overfitting settings. To evaluate the performance more realistically, we consider the two following setups for fine-tuning: (1) Over-training. Following the previous works [11, 23, 13], we fine-tune 15,000 steps on Pokemon datasets, and 150,000 steps on MS-COCO and Flickr (with only $/2500$ 2500 dataset size). (2) Real-world training. Considering that trainers typically do not train for such high steps, we recalibrate the steps based on the training steps/dataset size ratio (approximately 20) of official fine-tuning scripts on Huggingface [19]. Thus, we train 7,500 steps, 50,000 steps and 200,000 steps for the Pokémon, MS-COCO and Flickr datasets, respectively. Additionally, we employ the default data augmentation (Random-Crop and Random-Flip [20]) in training codes [20] to simulate real-world scenarios.
@@
Evaluation metrics. We follow the widely used metrics of previous works [3, 4, 11, 23, 13], including ASR (i.e., the accuracy of membership inference), AUC and the True Positive Rate (TPR) when the False Positive Rate (FPR) is 1% (i.e., TPR@1%FPR).
@@ 4.3 Performance on Various Training Steps
From Tab. 1 and Tab. 2, we find that the training steps greatly influence the effectiveness of membership inference. All membership inference methods tend to exhibit satisfactory performance when the model is trained for an excessive number of steps that conflicts with real-world scenarios. Therefore, we emphasize that the effectiveness trajectory of membership inference across varying training steps should also be utilized to evaluate different methods. Better membership inference methods should reveal membership information earlier as training progresses.
@@ 4.4 Ablation Study
Figure 3: Performance of $\text{CLiD}_{th}$ under various Monte Carlo sampling numbers (i.e., query count). The legend labels are sorted in ascending order of AUC values and SecMI. $M/N=0$ indicates that the corresponding term is not utilized in Eq. (16).
@@
Monte Carlo sampling numbers. In Fig. 3, we observe that when setting $M=N$, the performance improves as the number of Monte Carlo sampling increases. And the performance is improved slightly when $M,N>3$. Hence, we set $M,N=3$ to ensure the balance between a low query count and satisfied performance. Moreover, the results of "M=1, N=1" and "SecMI" also demonstrate that $\text{CLiD}_{th}$ outperforms previous works even with a fewer query count.
@@ 4.5 Resistance to Defense
Since data augmentation is commonly used in training and can mitigate the effectiveness of membership inference [11], we use it to evaluate the performance of methods under defense. As the baseline methods already exhibit weak performance under real-world training setting, we opt not to incorporate additional data augmentation. Instead, we remove the default data augmentation from training scripts [20] to observe the effectiveness change of different methods. We fine-tune Stable Diffusion models for 50,000 steps with MS-COCO, report the metrics, and calculate the metrics changes in Tab. 3. We observe that the effectiveness of all membership inference methods declines after data augmentation is introduced during training. Note that PFAMI [13] exhibits the highest sensitivity to data augmentation since it infers membership by probability fluctuation after images are perturbed, which also explains its significant performance decline between Tab. 1 and Tab. 2. Compared to the baselines, our method exhibits the smallest decrease, which indicates its stronger resistance to data augmentation.
Adaptive defense. We further consider adaptive defense: assuming the trainers are aware of our methods and perturb the text of image-text datasets before training. We consider the following adaptive defense methods: (1) rephrasing the original text  [^3] We utilize ChatGPT-3.5 with the following prompt: ”Please rewrite the following sentences while keeping the key semantics.”, (2) randomly deleting 10%, 30% words in text, and (3) shuffling 50% of the image-text mappings in the dataset. In Tab. 5, we observe that except for shuffling, the other adaptive defense methods have almost no effect on $\text{CLiD}_{th}$. And shuffling damages the model utility (too high FID values), rendering this defense meaningless.
@@
ASR AUC TPR@1%FPR ASR ($\Delta$) AUC ($\Delta$) TPR@1%FPR ($\Delta$)
@@ 4.6 Performance on Pre-trained Models
In this part, we evaluate the effectiveness of membership inference methods for the pre-training setting. To ensure the distribution consistency between the member and hold-out set, we respectively select $2500$ samples from the LAION-Aesthetics v2 5+ and LAION-2B MultiTranslated [45] as member/hold-out set following [12], and conduct membership inference on Stable Diffusion v1-5 [40]. As shown in Tab. 5, our method consistently outperforms the baselines across all three metrics.
@@ 5 Related Works
Membership inference on diffusion models. Recently, several membership inference methods on diffusion models have been proposed. In the grey-box or white-box setting, Carlini et al. [4] firstly conduct membership inference on unconditional diffusion models by conducting LiRA (Likelihood Ratio Attack) [3], with the requirement of training multiple shadow models. Matsumoto et al. [32] make the first step by utilizing diffusion loss to conduct query-based membership inference. Some works [11, 23] leverage the DDIM [48] deterministic forward process [22] to access the posterior estimation errors of diffusion models. And Fu et al. [13] leverage the probability fluctuations by perturbing image samples. Few works consider the black-box settings [54, 35]. However, these studies either assume partial knowledge of member set data [54] or assume extensive fine-tuning steps ($100\sim 500$ epochs) [35], both of which do not align with real-world scenarios.
@@ 6 Conclusion
To extensively validate the effectiveness of Assumption 3.1, we utilize additional metrics as the distances metric $D$ in Eq. (8), including Wasserstein Distance [55], Kernel MMD (Maximum Mean Discrepancy) [55] and 1-Nearest Neighbor Classifier (1-NN) [31], in addition to FID [15]. As observed in Fig. A.1, regardless of the metric used for $D$, Assumption 3.1 consistently holds, thereby confirming the broad applicability of Conditional Overfitting phenomenon.
@@ D.1 Monte Carlo Sampling
In Fig. D.1, we observe that the single Monte Carlo estimation achieves optimal accuracy when $t_{i}\in[400,500]$. Similar results are shown in [28]. Therefore, consistent with [28], we sample at intervals of $10$ centered around the timestep $450$. For example, if the estimation number (i.e., $M,N$ in Eq (14)) is $3$, we use the time list of $[440,450,460]$. Note that [4, 11] indicate that for DDPM of Cifar-10 [25], the best estimation timestep is around $100$. This difference may arise from the different signal-to-noise ratios of images with various resolution [18]. This finding suggests that the Monte Carlo sampling timestep should be designed differently for diffusion models of different scales.
@@ D.2 Reduction Methods
In implementation, we actually diversely reduce the condition $\mathbf{c}$ to $\mathbf{c}^{*}$ and calculate $p_{\theta}(\mathbf{x}|\mathbf{c}^{*})$ to approximate $p_{\theta}(\mathbf{x})$. In this part, we evaluate the effectiveness of different reduction methods. We consider three methods in Sec. 3.5: (1) Simply Clipping. We simply use the first, middle, and last thirds of the sentences as text inputs. (2) Gaussian Noise. We add Gaussian noises with the scales of $50\%,70\%,90\%$ to the overall text embeddings. (3) Importance Clipping. We calculate the importance of words in the text [^4] https://github.com/ma-labo/PromptCharm [49, 51] and replace them with “pad” tokens in descending order by varying proportions of $30\%,50\%,70\%$. For all three methods, we additionally use the null text as a $\mathbf{c}^{*}$. The experiments are conducted on the real-world training setting with MS-COCO dataset. And we also employ null text solely to compute Eq. (14) without reduction methods for comparison.
@@ Appendix E Compute Overhead and Resources
Computational Overhead. As a query-based member inference method, the computational efficiency of our method primarily depends on the number of queries. A lower query count signifies a more efficient member inference method. Our method significantly outperforms the baselines when the query count are about the same (such as SecMI and PFAMI in Sec. 4.2). Furthermore, even with a much lower query count such as $M=1,N=1(Q=5)$ (Fig. 3), our method exhibits a noticeable improvement compared to the baselines.
Compute Resources. Our experiments are divided into two main parts: training (fine-tuning) and inference, both conducted on a single RTX A6000 GPU. The time of execution in the training phase depends on the training steps. For example, we perform $7,500$, $50,000$, and $200,000$ steps for Pokemon [27], MS-COCO [29] and Flickr [57] dataset, which take about 2 hours, 12 hours, and 48 hours, respectively. The time of execution in inference time depends on the methods’ query count. For example, with the query count of $15$, our membership inference method on a dataset of size $/2500$ 2500 takes approximately $80$ minutes per run for all data points. Typically, we perform this inference once on the shadow model and once on the target model, resulting in a total time cost of $160$ minutes.
@@ Appendix G Limitation
Despite the significant improvements in membership inference of text-to-image diffusion models on various data distributions and data sizes, this work still has limitations. First, due to the limited availability of open-sourced pre-trained weights of text-to-image diffusion models, evaluations under the pre-training setting are not sufficiently comprehensive. Second, the Conditional Overfitting phenomenon discussed in this paper is only explored within the domain of text-to-image diffusion models. From an insight perspective, its presence may not be such limited. Future efforts could explore this in conditional diffusion models across different modalities such as text-to-video, and further investigate more universally applicable membership inference methods.